# Orbital Stability of Plane Waves in the Klein-Gordon Equation against Localized Perturbations

### v1

@@ Abstract.
We investigate the stability and long-term behavior of spatially periodic plane waves in the complex Klein-Gordon equation under localized perturbations. Such perturbations render the wave neither localized nor periodic, placing its stability analysis outside the scope of the classical orbital stability theory for Hamiltonian systems developed by Grillakis, Shatah, and Strauss. Inspired by Zhidkov’s work on the stability of time-periodic, spatially homogeneous states in the nonlinear Schrödinger equation, we develop an alternative method that relies on an amplitude-phase decomposition and leverages conserved quantities tailored to the perturbation equation. We establish an orbital stability result of plane waves that is locally uniform in space, accommodating $L^{2}$-localized perturbations as well as nonlocalized phase modulations. In certain regimes, our method even allows for unbounded modulations. Our result is sharp in the sense that it holds up to the spectral stability boundary.
@@ 1. Introduction
The Klein-Gordon equation arises in a wide array of physical contexts, including nonlinear wave propagation, superconductivity, and quantum field dynamics; see, for instance, the classical textbooks [Lee81, RS75, Whi99] and references therein. In many applications, the nonlinearity takes a specific form, such as the power law $f(\nu)=\smash{1\pm\nu^{p}}$ with $p\in\mathbb{N}$. In this work, however, we consider a general class of nonlinearities and impose no specific structural assumptions on $f$, beyond its ${\mathcal{C}}^{2}$-regularity.
@@
$H(\xi,\zeta,\mu,\nu)=\frac{1}{2}\int_{\mathcal{I}}\xi^{\prime}(x)^{2}+\zeta^{\prime}(x)^{2}+\mu(x)^{2}+\nu(x)^{2}+F\big{(}\xi(x)^{2}+\zeta(x)^{2}\big{)}\,\mathrm{d}x,$
@@
Conservation laws play a crucial role in understanding the long-term dynamics of Hamiltonian systems. Their use in proving global existence or finite-time blow-up of solutions is classical; see, for example, [CH98] for applications to the Klein–Gordon equation. Conserved quantities also form the backbone of the orbital stability theory of traveling (or standing) wave solutions. The primary approach to proving orbital stability of traveling waves is to characterize them as constrained minimizers of the Hamiltonian. This variational approach, dating back to ideas of Boussinesq, was first introduced in the nonlinear stability analysis of solitary waves in the Korteweg-de Vries equation by Benjamin [Ben72]. Its first application to the Klein-Gordon equation, resulting in orbital stability of standing waves, is due to Shatah [Sha83]. Building on the work of Benjamin and others, the method was formalized by Grillakis, Shatah, and Straus, culminating in a comprehensive orbital stability theory [GSS87, GSS90] for traveling waves in general Hamiltonian systems with symmetries. A thorough introduction on the method and further references can be found in [KP13].
Although the theory of Grillakis, Shatah, and Strauss provides an adequate framework for the stability analysis of solitary waves, its application to periodic waves faces limitations. Specifically, the variational approach is heavily dependent on conserved quantities (such as the Hamiltonian), which require that the perturbed wave is either localized or periodic. As a result, the class of admissible perturbations for a periodic wave is restricted to those with the same period (co-periodic or harmonic perturbations) or those whose period is a multiple of that of the wave (subharmonic perturbations). In the Klein-Gordon equation (1.1), orbital stability results of periodic waves against harmonic perturbations have been obtained for several classes of nonlinearities; see [CP23, NPF08, Pal22]. For an overview of co-periodic stability results in other Hamiltonian systems, as well as a general characterization of the Grillakis-Shatah-Strauss stability criteria in the periodic context, we refer to [BGNR13, BGMR16].
As outlined in [BGNR13, BGMR16, GH07], the nonlinear stability of periodic waves with respect to localized perturbations in Hamiltonian systems remains a largely unsolved problem, standing in sharp contrast with the dissipative case; see Remark 1. To the authors’ best knowledge, the only known nonlinear stability result to date for this class of perturbations concerns the orbital stability of plane waves in the one-dimensional nonlinear Schrödinger (NLS) equation
@@
to (1.6). Its orbital stability with respect to localized perturbations was established by Zhidkov [Zhi01]. The result holds locally uniformly in space and is derived via variational methods: by taking a suitable formal linear combination of the classical conservation laws for mass and energy, one obtains a well-defined conserved quantity, which serves to close the nonlinear argument. As observed in [GH07], this also implies orbital stability of the associated plane-wave solution under localized perturbations.
@@
The nonlinear stability theory of periodic waves in dissipative systems with respect to localized perturbations has advanced significantly over the last decades. The first nonlinear stability result was obtained for plane-wave solutions in the real Ginzburg-Landau equation [CEE92] by exploiting its gauge invariance. The use of mode filters in Bloch frequency domain [Sch96, Sch98], combined with the modulational ansatz proposed in [DSSS09], led to a breakthrough in the nonlinear stability analysis of periodic waves in large classes of dissipative systems such as reaction-diffusion models [JZ11, JNRZ13, SSSU12] and systems of viscous conservation laws [JZ10, JNRZ14]. Based on these developments, we believe that the orbital stability analysis of plane waves in the complex Klein-Gordon equation presented in this paper offers a promising step toward developing an orbital stability theory for more general periodic waves in Hamiltonian systems.
@@ 1.1. Goal, challenges, and strategy
Our aim is to develop a variational framework, in the spirit of [Zhi01], for establishing orbital stability of plane waves in the complex Klein-Gordon equation (1.1) against localized perturbations. In contrast to the NLS equation, the Klein-Gordon equation is not completely integrable, which significantly limits the number of (formally) conserved quantities that could be used in a nonlinear stability argument. In addition, unlike the NLS equation, where Galilean invariance enables a reduction to a spatially homogeneous solution, no such reduction is available for the Klein-Gordon equation (1.1).
To address these difficulties, we adopt an alternative reduction inspired by the spectral stability analysis [DHSS15] of spatially periodic waves in (1.1). Specifically, we introduce the new coordinate
@@
If the second variation ′′$\mathcal{H}^{\prime\prime}(\mathbf{w}_{*})$ were positive definite, then the energy acts as a Lyapunov functional and $\|\mathbf{z}(t)\|^{2}$ can be bounded in terms of $E(\mathbf{z}(t))=E(\mathbf{z}_{0})$, implying nonlinear stability. However, this is obstructed by the gauge symmetry of (1.1) which forces 00 to lie in the spectrum of ′′$\mathcal{H}^{\prime\prime}(\mathbf{w}_{*})$. We emphasize that, since ′′$\mathcal{H}^{\prime\prime}(\mathbf{w}_{*})$ has constant coefficients, there is also no finite-codimensional subspace of $L^{2}({\mathbb{R}})$ on which ′′$\mathcal{H}^{\prime\prime}(\mathbf{w}_{*})$ is positive definite, which is a key assumption in the classical Grillakis-Shatah-Strauss stability framework.
@@
ensuring that ′′$\mathcal{H}^{\prime\prime}(\mathbf{w}_{*})$ is positive semi-definite, is satisfied. All in all, the obtained control on $\rho(t)$ and on derivatives of $\phi(t)$ is sufficient to close a nonlinear iteration argument, yielding an orbital stability result which is locally uniform in space.
@@
where the wave speed $s$, wave number $k$, and temporal frequency $\omega$ are real parameters and the wave profile $\psi_{0}\colon{\mathbb{R}}\to{\mathbb{R}}$ connects nonvanishing asymptotic end states $\psi_{\pm}=\lim_{\xi\to\pm\infty}\psi(\xi)\neq 0$. When the wave profile is strictly monotone, i.e. $\psi_{0}^{\prime}(\xi)\neq 0$ for all $\xi\in{\mathbb{R}}$, the use of variational methods to prove orbital stability is well-established in one-dimensional Hamiltonian systems. We refer to [BGSS08, GZ09, HPW82, IK87, Zhi87a, Zhi87b, Zhi01] for results in the (generalized) Korteweg-de Vries, Klein-Gordon, NLS, and nonlinear wave equations. A further extension to solitonic bubble solutions of the form $\psi(x,t)=\psi_{0}(x-st)\mathrm{e}^{\mathrm{i}\phi_{0}(x-st)}$, with $\psi_{0}$ even and $(\psi_{0}(\xi),\phi_{0}^{\prime}(\xi))\to(\psi_{\infty},0)$ as $\xi\to\pm\infty$ can be found in [Lin02].
@@ 1.2. Main result
We formulate our main result concerning the orbital stability of plane waves in the complex Klein-Gordon equation under localized perturbations. Since our analysis only requires control over the $L^{2}$-norm of the spatial and temporal derivatives of the phase variable, we are able to establish orbital stability under nonlocalized initial phase modulations. Thus, we consider initial data of the form
$\begin{pmatrix}u(x,0)\\ \partial_{t}u(x,0)\end{pmatrix}=a\mathrm{e}^{\mathrm{i}kx-\mathrm{i}\theta_{\infty}(x)}\begin{pmatrix}1\\ \mathrm{i}\omega\end{pmatrix}+\begin{pmatrix}w_{0}(x)\\ v_{0}(x)\end{pmatrix}$ (1.13)
where $\theta_{\infty}\colon{\mathbb{R}}\to{\mathbb{R}}$ is a continuously differentiable phase function with $\theta_{\infty}^{\prime}\in H^{1}({\mathbb{R}})$, and $w_{0}\in H^{1}({\mathbb{R}})$ and $v_{0}\in L^{2}({\mathbb{R}})$ are localized perturbations. The sign of $f(a^{2})$ plays an important role in our analysis and determines whether unbounded phase modulations are admissible or not. This gives rise to the following definition.
Definition 3.
Let $a>0$. We say that $\theta_{\infty}\in{\mathcal{C}}^{1}({\mathbb{R}},{\mathbb{R}})$ is $a$-admissible if $\theta_{\infty}^{\prime}\in H^{1}({\mathbb{R}})$ and if one of the following two scenarios holds.
  1. (i)
    $f(a^{2})>0$,
  2. (ii)
    $f(a^{2})\leq 0$ and $\theta_{\infty}^{\prime}\in L^{1}({\mathbb{R}})$.
If $\theta_{\infty}$ is $a$-admissible, we then write
[…]
Thus, unbounded initial phase modulations $\theta_{\infty}\in{\mathcal{C}}^{1}({\mathbb{R}})$ in (1.13) are $a$-admissible if $f(a^{2})>0$ and $\theta_{\infty}^{\prime}\in H^{1}({\mathbb{R}})$. In contrast, if $f(a^{2})\leq 0$, then only bounded phase modulations $\theta_{\infty}$ are $a$-admissible. We refer to Remarks 5 and 6 for further details.
@@
Let $(a,k,\omega)\in(0,\infty)\times{\mathbb{R}}^{2}$ with $(k,\omega)\neq(0,0)$. Assume that the existence condition (1.3) and the spectral condition (1.12) are satisfied. Fix $R>0$. Then, there exist constants $C,\delta>0$ such that, whenever $\theta_{\infty}\in{\mathcal{C}}^{1}({\mathbb{R}},{\mathbb{R}})$ is $a$-admissible and satisfies
$E_{\infty}\leq\delta,$ (1.14)
the following holds.
If $(w_{0},v_{0})\in H^{2}({\mathbb{R}})\times H^{1}({\mathbb{R}})$ obeys
$E_{0}:=\lVert w_{0}\rVert_{H^{1}({\mathbb{R}})}+\lVert v_{0}\rVert_{L^{2}({\mathbb{R}})}\leq\delta,$ (1.15)
then there exists a unique global classical solution
@@
    (Local orbital stability) For every $x_{*}\in{\mathbb{R}}$ and $t\geq 0$ there exists a constant $\gamma=\gamma(x_{*},t)\in{\mathbb{R}}$ such that
    $\lVert u(t)-a\mathrm{e}^{\mathrm{i}k(\cdot)+\mathrm{i}\omega t+\mathrm{i}\gamma}\rVert_{H^{1}(x_{*}-R,x_{*}+R)}\leq C(E_{0}+E_{\infty}).$ (1.17)
@@
    $u(x,t)=a\mathrm{e}^{\mathrm{i}kx+\mathrm{i}\omega t-\mathrm{i}\theta_{\infty}(x)}\,\mathrm{e}^{\varrho(x,t)+\mathrm{i}\vartheta(x,t)}$ (1.18)
@@
    $\begin{split}\lVert\varrho_{t}(t)\rVert_{L^{2}({\mathbb{R}})}+\lVert\vartheta_{t}(t)\rVert_{L^{2}({\mathbb{R}})}+\lVert\varrho(t)\rVert_{H^{1}({\mathbb{R}})}+\lVert\vartheta_{x}(t)\rVert_{L^{2}({\mathbb{R}})}&\leq C(E_{0}+E_{\infty}),\\ \lVert\vartheta(t)\rVert_{L^{2}({\mathbb{R}})}&\leq C(E_{0}+E_{\infty})(t+1)\end{split}$ (1.19)
@@
In the regime $f(a^{2})>0$, Theorem 4 allows for unbounded phase modulations $\theta_{\infty}$. For instance, there exists a constant $C_{0}>0$ such that the unbounded smooth function $\theta_{\infty}(x)=\smash{(1+\epsilon^{4}x^{2})^{\frac{1}{8}}}$ satisfies $\|\theta_{\infty}^{\prime}\|_{H^{1}({\mathbb{R}})}\leq C_{0}\epsilon$ for all $\epsilon\in(0,1)$. Thus, for any fixed $\delta,K>0$, the condition (1.14) can be met by taking $\epsilon\in(0,1)$ sufficiently small. In contrast, when $f(a^{2})\leq 0$, we impose the additional requirement $\theta_{\infty}^{\prime}\in L^{1}({\mathbb{R}})$. This condition guarantees that $\theta_{\infty}$ is bounded, its asymptotic limits $\Theta_{\pm\infty}=\lim_{x\to\pm\infty}\theta_{\infty}(x)$ exist, and the phase offset $\Theta_{\infty}-\Theta_{-\infty}$ is bounded by $\|\theta_{\infty}^{\prime}\|_{L^{1}({\mathbb{R}})}$. As a result, the phase offset must remain small when $f(a^{2})\leq 0$. As an illustrative example, consider
$\theta_{\infty}(x)=kx_{-}\frac{1-\tanh(x)}{2}+kx_{+}\frac{1+\tanh(x)}{2}$
[…]

### v2

# arXiv:2506.06029v2[math.AP] 01 Mar 2026

@@ Abstract.
We investigate the stability and long-term behavior of spatially periodic plane waves in the complex Klein-Gordon equation under localized perturbations. Such perturbations render the wave neither localized nor periodic, placing its stability analysis outside the scope of the classical orbital stability theory for Hamiltonian systems developed by Grillakis, Shatah, and Strauss. Inspired by Zhidkov’s work on the stability of time-periodic, spatially homogeneous states in the nonlinear Schrödinger equation, we develop an alternative method that relies on an amplitude-phase decomposition and leverages conserved quantities tailored to the perturbation equation. We establish an orbital stability result of plane waves that is locally uniform in space, accommodating $L^{2}$-localized perturbations as well as unbounded phase modulations. Our result is sharp in the sense that it holds up to the spectral stability boundary.
@@ 1. Introduction
The Klein-Gordon equation arises in a wide array of physical contexts, including nonlinear wave propagation, superconductivity, and quantum field dynamics; see, for instance, the classical textbooks [27, 31, 37] and references therein. In many applications, the nonlinearity takes a specific form, such as the power law $f(\nu)=\smash{1\pm\nu^{p}}$ with $p\in\mathbb{N}$. In this work, however, we consider a general class of nonlinearities and impose no specific structural assumptions on $f$, beyond its ${\mathcal{C}}^{2}$-regularity.
@@
$H(\xi,\zeta,\mu,\nu)=\frac{1}{2}\int_{\mathcal{I}}\xi^{\prime}(x)^{2}+\zeta^{\prime}(x)^{2}+\mu(x)^{2}+\nu(x)^{2}+F\big(\xi(x)^{2}+\zeta(x)^{2}\,\mathrm{d}x,$
@@
Conservation laws play a crucial role in understanding the long-term dynamics of Hamiltonian systems. Their use in proving global existence or finite-time blow-up of solutions is classical; see, for example, [7] for applications to the Klein–Gordon equation. Conserved quantities also form the backbone of the orbital stability theory of traveling (or standing) wave solutions. The primary approach to proving orbital stability of traveling waves is to characterize them as constrained minimizers of the Hamiltonian. This variational approach, dating back to ideas of Boussinesq, was first introduced in the nonlinear stability analysis of solitary waves in the Korteweg-de Vries equation by Benjamin [2]. Its first application to the Klein-Gordon equation, resulting in orbital stability of standing waves, is due to Shatah [36]. Building on the work of Benjamin and others, the method was formalized by Grillakis, Shatah, and Straus, culminating in a comprehensive orbital stability theory [16, 17] for traveling waves in general Hamiltonian systems with symmetries. A thorough introduction on the method and further references can be found in [25].
Although the theory of Grillakis, Shatah, and Strauss provides an adequate framework for the stability analysis of solitary waves, its application to periodic waves faces limitations. Specifically, the variational approach is heavily dependent on conserved quantities (such as the Hamiltonian), which require that the perturbed wave is either localized or periodic. As a result, the class of admissible perturbations for a periodic wave is restricted to those with the same period (co-periodic or harmonic perturbations) or those whose period is a multiple of that of the wave (subharmonic perturbations). In the Klein-Gordon equation (1.1), orbital stability results of periodic waves against harmonic perturbations have been obtained for several classes of nonlinearities; see [8, 29, 30]. For an overview of co-periodic stability results in other Hamiltonian systems, as well as a general characterization of the Grillakis-Shatah-Strauss stability criteria in the periodic context, we refer to [4, 3].
As outlined in [4, 3, 14], the nonlinear stability of periodic waves with respect to localized perturbations in Hamiltonian systems remains a largely unsolved problem, standing in sharp contrast with the dissipative case; see Remark 1. To the authors’ best knowledge, the only known nonlinear stability result to date for this class of perturbations concerns the orbital stability of plane waves in the one-dimensional nonlinear Schrödinger (NLS) equation
@@
to (1.6). Its orbital stability with respect to localized perturbations was established by Zhidkov [40]. The result holds locally uniformly in space and is derived via variational methods: by taking a suitable formal linear combination of the classical conservation laws for mass and energy, one obtains a well-defined conserved quantity, which serves to close the nonlinear argument. As observed in [14], this also implies orbital stability of the associated plane-wave solution under localized perturbations.
@@
The nonlinear stability theory of periodic waves in dissipative systems with respect to localized perturbations has advanced significantly over the last decades. The first nonlinear stability result was obtained for plane-wave solutions in the real Ginzburg-Landau equation [9] by exploiting its gauge invariance. The use of mode filters in Bloch frequency domain [34, 35], combined with the modulational ansatz proposed in [12], led to a breakthrough in the nonlinear stability analysis of periodic waves in large classes of dissipative systems such as reaction-diffusion models [24, 21, 33] and systems of viscous conservation laws [23, 22]. Based on these developments, we believe that the orbital stability analysis of plane waves in the complex Klein-Gordon equation presented in this paper offers a promising step toward developing an orbital stability theory for more general periodic waves in Hamiltonian systems.
@@ 1.1. Goal, challenges, and strategy
Our aim is to develop a variational framework, in the spirit of [40], for establishing orbital stability of plane waves in the complex Klein-Gordon equation (1.1) against localized perturbations. In contrast to the NLS equation where Galilean invariance enables a reduction to a spatially homogeneous solution to the NLS equation itself, no such reduction is available for the Klein-Gordon equation (1.1).
To address this difficulty, we adopt an alternative reduction inspired by the spectral stability analysis [11] of spatially periodic waves in (1.1). Specifically, we introduce the new coordinate
@@
If the second variation $\mathcal{H}^{\prime\prime}(\mathbf{w}_{*})$ were positive definite, then the energy acts as a Lyapunov functional and $\|\mathbf{z}(t)\|^{2}$ can be bounded in terms of $E(\mathbf{z}(t))=E(\mathbf{z}_{0})$, implying nonlinear stability. However, this is obstructed by the gauge symmetry of (1.1) which forces 0 to lie in the spectrum of $\mathcal{H}^{\prime\prime}(\mathbf{w}_{*})$. We emphasize that, since $\mathcal{H}^{\prime\prime}(\mathbf{w}_{*})$ has constant coefficients, there is also no finite-codimensional subspace of $L^{2}({\mathbb{R}})$ on which $\mathcal{H}^{\prime\prime}(\mathbf{w}_{*})$ is positive definite, which is a key assumption in the classical Grillakis-Shatah-Strauss stability framework.
@@
ensuring that $\mathcal{H}^{\prime\prime}(\mathbf{w}_{*})$ is positive semi-definite, is satisfied. All in all, the obtained control on $\rho(t)$ and on derivatives of $\phi(t)$ is sufficient to close a nonlinear iteration argument, yielding an orbital stability result which is locally uniform in space.
@@
where the wave speed $s$, wave number $k$, and temporal frequency $\omega$ are real parameters and the wave profile $\psi_{0}\colon{\mathbb{R}}\to{\mathbb{R}}$ connects nonvanishing asymptotic end states $\psi_{\pm}=\lim_{\xi\to\pm\infty}\psi(\xi)\neq 0$. When the wave profile is strictly monotone, i.e. $\psi_{0}^{\prime}(\xi)\neq 0$ for all $\xi\in{\mathbb{R}}$, the use of variational methods to prove orbital stability is well-established in one-dimensional Hamiltonian systems. We refer to [5, 15, 18, 20, 38, 39, 40] for results in the (generalized) Korteweg-de Vries, Klein-Gordon, NLS, and nonlinear wave equations. A further extension to solitonic bubble solutions of the form $\psi(x,t)=\psi_{0}(x-st)\mathrm{e}^{\mathrm{i}\phi_{0}(x-st)}$, with $\psi_{0}$ even and $(\psi_{0}(\xi),\phi_{0}^{\prime}(\xi))\to(\psi_{\infty},0)$ as $\xi\to\pm\infty$ can be found in [28].
@@ 1.2. Main result
We formulate our main result concerning the orbital stability of plane waves in the complex Klein-Gordon equation under localized perturbations. Since our analysis only requires control over the $L^{2}$-norm of the spatial and temporal derivatives of the phase variable, we are able to establish orbital stability under unbounded initial phase modulations; see Remark 4. Thus, we consider initial data of the form
$\begin{pmatrix}u(x,0)\\ \partial_{t}u(x,0)\end{pmatrix}=a\mathrm{e}^{\mathrm{i}kx+\mathrm{i}\theta_{\infty}(x)}\begin{pmatrix}1\\ \mathrm{i}\omega\end{pmatrix}+\begin{pmatrix}w_{0}(x)\\ v_{0}(x)\end{pmatrix}$ (1.13)
where $\theta_{\infty}\colon{\mathbb{R}}\to{\mathbb{R}}$ is a continuously differentiable phase function with $\theta_{\infty}^{\prime}\in H^{1}({\mathbb{R}})$, and $w_{0}\in H^{1}({\mathbb{R}})$ and $v_{0}\in L^{2}({\mathbb{R}})$ are localized perturbations.
@@
Let $(a,k,\omega)\in(0,\infty)\times{\mathbb{R}}^{2}$ with $(k,\omega)\neq(0,0)$. Assume that the existence condition (1.3) and the spectral condition (1.12) are satisfied. Then, there exist constants $C,\delta>0$ such that, whenever $\theta_{\infty}\in{\mathcal{C}}^{1}({\mathbb{R}},{\mathbb{R}})$ and $(w_{0},v_{0})\in H^{2}({\mathbb{R}})\times H^{1}({\mathbb{R}})$ satisfy $\theta_{\infty}^{\prime}\in H^{1}({\mathbb{R}})$ and
$E_{0}:=\lVert\theta_{\infty}^{\prime}\rVert_{L^{2}({\mathbb{R}})}+\lVert w_{0}\rVert_{H^{1}({\mathbb{R}})}+\lVert v_{0}\rVert_{L^{2}({\mathbb{R}})}\leq\delta,$ (1.14)
there exists a unique global classical solution
@@
    (Local orbital stability) For every $R>0$, $x_{*}\in{\mathbb{R}}$, and $t\geq 0$ there exists a constant $\gamma=\gamma(x_{*},t)\in{\mathbb{R}}$ such that
    $\lVert u(t)-a\mathrm{e}^{\mathrm{i}k(\cdot)+\mathrm{i}\omega t+\mathrm{i}\gamma}\rVert_{H^{1}(x_{*}-R,x_{*}+R)}\leq C\sqrt{R}E_{0}.$ (1.16)
@@
    $u(x,t)=a\mathrm{e}^{\mathrm{i}kx+\mathrm{i}\omega t+\mathrm{i}\theta_{\infty}(x)}\,\mathrm{e}^{\varrho(x,t)+\mathrm{i}\vartheta(x,t)}$ (1.17)
@@
    $\begin{split}\lVert\varrho_{t}(t)\rVert_{L^{2}({\mathbb{R}})}+\lVert\vartheta_{t}(t)\rVert_{L^{2}({\mathbb{R}})}+\lVert\varrho(t)\rVert_{H^{1}({\mathbb{R}})}+\lVert\vartheta_{x}(t)\rVert_{L^{2}({\mathbb{R}})}&\leq CE_{0},\\ \lVert\vartheta(t)\rVert_{L^{2}({\mathbb{R}})}&\leq CE_{0}(t+1)\end{split}$ (1.18)
@@
As an illustrative example for the initial phase modulation in Theorem 3, consider
$\theta_{\infty}(x)=kx_{-}\frac{\tanh(x)-1}{2}-kx_{+}\frac{1+\tanh(x)}{2}$
with $k,x_{\pm}\in{\mathbb{R}}$. This function satisfies $\lVert\theta_{\infty}^{\prime}\rVert_{L^{2}({\mathbb{R}})}=|k(x_{+}-x_{-})|/\sqrt{3}$. Hence, for any fixed $\delta>0$, it obeys $\|\theta_{\infty}^{\prime}\|_{L^{2}({\mathbb{R}})}<\delta$ provided that the phase offset $k(x_{+}-x_{-})$ is sufficiently small. The associated modulated plane wave $\smash{a\mathrm{e}^{\mathrm{i}kx+\mathrm{i}\theta_{\infty}(x)}}$ connects to the phase-shifted waves $\smash{a\mathrm{e}^{\mathrm{i}k(x-x_{\pm})}}$ at $\pm\infty$. Theorem 3 also allows for unbounded initial phase modulations $\theta_{\infty}$. For instance, there exists a constant $C_{0}>0$ such that the unbounded smooth function $\theta_{\infty}(x)=\smash{(1+\epsilon^{4}x^{2})^{\frac{1}{8}}}$ satisfies $\|\theta_{\infty}^{\prime}\|_{L^{2}({\mathbb{R}})}\leq C_{0}\epsilon$ for all $\epsilon\in(0,1)$. Thus, for any fixed $\delta>0$, the condition (1.14) can be met by taking $\epsilon\in(0,1)$ sufficiently small.